OpenStudy (anonymous):

Can anybody help me on this Solve the equation for y. Then graph the equation. 5x – 4y = 12

OpenStudy (anonymous):

I would start by subtracting 5x from both sides, then divide both side by -4 and it should come out to be y=5/4x-3

OpenStudy (anonymous):

b= 12/-4 is not right

OpenStudy (anonymous):

where did b=12/-4 come into the equation?

OpenStudy (anonymous):

sorry I wrote it wrong

OpenStudy (anonymous):

this is what I had an got it wrong B. y = 4/5x + 3

OpenStudy (anonymous):

where dose the y=5/-4x-3 come from or is it y=5/4x-3

OpenStudy (anonymous):

5x – 4y = 12 5x – 4y-5x = 12-5x (subtract 5x from both sides) -4y=-5x+12 (simplify) -4/-4y=-5/-4x+12/-4(divided by -4 on both sides) y=5/4x-3 (simplify)

OpenStudy (anonymous):

oh ok thanks a lot